If the average monthly salary of two builders is £2210 and three times their difference is 540 find their salaries.
Answer by Theo(13342)

let x = the salary of the first builder.
let y = the salary of the second builder.

their average monthly salary is 2210.
this means that (x + y) /2 = 2210
multiply both sides of this equation by 2 and you get x + y = 4420

3 times their difference is 540.
this means that 3 * (x - y) = 540
divide both sides of this equation by 3 and you get x - y = 180

you have 2 equations that need to be solved simultaneously.

they are:

x + y = 4420
x - y = 180

add both sides of each equation together and you get:

2x = 4600

divide both sides of the equation by 2 and you get x = 2300

since x + y = 4420, this means that y must be equal to 4420 - 2300 = 2120

you have x = 2300 and y = 2120.

their average is (2300 + 2120) / 2 = 2210

3 times their difference is 3 * (2300 - 2120) = 3 * 180 = 540

that's your solution.

the salary of one of the builders is 2300 and the salary of the other builder is 2120.